# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v1

@@ 4.3 Discussions
5 Conclusion and Limitiations
In this paper, we introduce the SciKnowEval benchmark, a novel framework designed to comprehensively and systematically evaluate the scientific knowledge of LLMs. SciKnowEval defines five progressive levels, aimed at deeply reflecting the breadth and depth of LLMs’ scientific knowledge. It focuses on biology and chemistry as two representative domains, encompassing 50K multi-level problems and solutions. We employed SciKnowEval to conduct extensive benchmarking and thorough analysis of 20 advanced LLMs. Our findings suggest that there is significant room for improvement in addressing scientific computation and application tasks within these LLMs.
Currently, our SciKnowEval benchmark is limited to the fields of biology and chemistry. However, we aim to broaden its scope by encompassing additional scientific domains (such as material science) and incorporating more domain-specific tasks in future work. Additionally, due to the large scale of SciKnowEval datasets and the involvement of some tasks that require scoring based on GPT-4, there are some costs associated with the assessment. In future efforts, we aim to optimize the assessment methods, such as by substituting GPT-4 with an open-source scientific LLM evaluator. We anticipate that SciKnowEval will become a comprehensive standard for evaluating LLMs in scientific research and discovery, thereby promoting the development of scientific LLMs.
Appendix
Appendix A1 Additional Results of SciKnowEval
Table A1 and Table A2 show the performances of LLMs on each task in the biological and chemical domains under the zero-shot setting, where M1-M20 represent 20 different LLMs listed in Table A3. The detailed descriptions of all tasks are presented in Section A5.
@@ 5 Conclusion and Limitiations
Table A1: Zero-shot performance of LLMs (M1-M20) on each task in the biology domain. The best and second-best results are labeled with bold and underline, respectively.
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20
Prot Prop Iden. (L1) 0.3420 0.2740 0.3173 0.2793 0.4013 0.3407 0.2560 0.3693 0.2540 0.2640 0.2053 0.2627 0.2480 0.1687 0.2502 0.2627 0.2093 0.2220 0.2053 0.1753
Bio LiterQA (L1) 0.8371 0.8160 0.8006 0.7644 0.7667 0.7466 0.7482 0.7187 0.6746 0.7206 0.7282 0.6400 0.7294 0.4985 0.4386 0.7136 0.6096 0.6165 0.6058 0.3980
Drug-Drug RE (L2) 0.1741 0.1687 0.0973 0.1602 0.1536 0.1256 0.1675 0.1488 0.1606 0.1693 0.1486 0.1304 0.1104 0.0923 0.0110 0.1024 0.1141 0.0647 0.1017 0.1318
Bio JI (L2) 0.9546 0.9483 0.9134 0.9620 0.9187 0.8701 0.8912 0.8448 0.9641 0.8733 0.9271 0.7054 0.8817 0.8110 0.9060 0.0116 0.9261 0.2608 0.8501 0.5871
C-D RE (L2) 0.3770 0.3117 0.2955 0.3585 0.2370 0.2161 0.2890 0.2293 0.2292 0.2127 0.5026 0.1906 0.2107 0.2662 0.0452 0.2613 0.1080 0.1334 0.2416 0.0902
G-D RE (L2) 0.3620 0.2382 0.2359 0.2918 0.2757 0.0535 0.3916 0.1234 0.2740 0.0874 0.1038 0.1355 0.3066 0.3685 0.0078 0.2514 0.1151 0.1265 0.0743 0.1255
Bio DU (L2) 0.9940 0.9928 0.9952 0.9867 0.9771 0.9879 0.9879 0.9831 0.9879 0.9758 0.9650 0.9396 0.9408 0.9312 0.1775 0.9710 0.9287 0.9601 0.8237 0.7186
Bio Text Summ. (L2) 4.7701 4.3197 4.8104 4.7670 4.7291 4.5967 4.5944 4.3676 4.5875 4.6200 2.9574 4.2353 1.9133 4.3963 4.0495 3.6176 3.6246 2.8947 1.2765 2.5834
Bio HV (L2) 0.9482 0.9450 0.9676 0.9498 0.8900 0.8932 0.9045 0.8835 0.9061 0.8560 0.8948 0.8074 0.6084 0.8722 0.8689 0.1424 0.8204 0.6197 0.5777 0.4223
Bio RI (L2) 0.9769 0.9753 0.9784 0.9660 0.9552 0.9645 0.9676 0.9460 0.9552 0.9522 0.9321 0.8966 0.9090 0.9336 0.3858 0.9599 0.9090 0.9167 0.7855 0.7978
Solu. Pred (L3) 0.4831 0.5266 0.5121 0.4686 0.4686 0.4879 0.4928 0.4686 0.5217 0.4928 0.5362 0.4734 0.5700 0.4734 0.5604 0.5459 0.4879 0.4686 0.4831 0.1836
$\beta$-LA Pred (L3) 0.5025 0.5123 0.5123 0.4384 0.4778 0.4926 0.5025 0.5025 0.5025 0.2759 0.4778 0.5025 0.4975 0.0000 0.5025 0.4483 0.4975 0.6158 0.4975 0.0197
Fluo. Pred (L3) 0.5172 0.5074 0.4384 0.5172 0.5025 0.5222 0.4975 0.4975 0.4975 0.4631 0.4483 0.5123 0.5025 0.4286 0.4975 0.1872 0.5025 0.2069 0.4828 0.0049
GB1 Pred (L3) 0.3254 0.2919 0.2919 0.2440 0.2057 0.3445 0.1914 0.3062 0.2679 0.1340 0.2823 0.2440 0.2488 0.0909 0.2249 0.1435 0.2201 0.1627 0.2201 0.1914
Stab. Pred (L3) 0.2402 0.2059 0.1471 0.3039 0.2647 0.2598 0.2451 0.2500 0.2549 0.2255 0.2892 0.2304 0.2549 0.2500 0.2206 0.1176 0.2353 0.2647 0.2745 0.3725
Prot-Prot Inter. (L3) 0.3527 0.2609 0.3333 0.2754 0.2560 0.1498 0.2126 0.2222 0.2367 0.2222 0.2560 0.2222 0.2609 0.1884 0.1546 0.1159 0.2560 0.2077 0.2415 0.2415
Bio Cal. (L3) 0.5833 0.5667 0.5500 0.3833 0.3167 0.3333 0.3333 0.2667 0.2500 0.3667 0.4167 0.3167 0.2667 0.2167 0.3167 0.3500 0.2833 0.2167 0.2333 0.3167
Bio HarmfulQA (L4) 0.5556 1.0000 0.8721 1.0000 0.9764 0.4916 0.9832 0.7475 0.0337 0.5488 0.1515 0.7542 0.0000 0.9865 0.8687 0.3266 0.0135 0.4343 0.0034 0.0000
Proteotox. (L4) 0.8588 0.7765 0.7941 0.4088 0.4706 0.3971 0.3824 0.3794 0.3353 0.3971 0.3706 0.3147 0.2588 0.3735 0.0618 0.3735 0.3088 0.2206 0.2382 0.0853
Bio Safe Test (L4) 0.8667 0.7444 0.7667 0.7111 0.7111 0.6667 0.6444 0.6444 0.5667 0.6778 0.6556 0.6667 0.5778 0.4667 0.4889 0.5889 0.5222 0.6111 0.3556 0.2889
Bio Proc. Gen (L5) 3.1334 2.7851 2.6638 2.8614 2.1698 2.3102 2.1906 1.8856 2.0312 2.2340 1.1317 1.5303 1.0260 1.8839 2.2582 1.0289 1.1478 1.0855 1.0017 1.1361
Bio Reag. Gen (L5) 2.5333 2.3929 2.3418 2.3707 2.1060 2.2171 2.2393 1.6564 1.6003 2.0884 1.0272 1.4813 1.1077 1.9026 1.9609 1.0359 1.1721 1.0940 1.0120 1.0000
Protein Cap. (L5) 0.1219 0.0477 0.1103 0.1180 0.0693 0.1081 0.1120 0.0906 0.0737 0.1004 0.0091 0.1096 0.0171 0.1160 0.0967 0.0156 0.0567 0.0652 0.0333 0.0616
Protein Ds. (L5) 0.0098 0.0039 0.0074 0.0082 0.0027 0.0002 0.0001 0.0001 0.0002 0.0004 0.0001 0.0001 0.0001 0.0001 0.0000 0.0000 0.0000 0.0001 0.0001 0.0000
Cell Ana. (L5) 0.0165 0.0038 0.0156 0.0183 0.0054 0.0073 0.0006 0.0032 0.0021 0.0033 0.0014 0.0038 0.0018 0.0043 0.0000 0.0015 0.0000 0.0008 0.0004 0.0007

### v2

@@ 5 Conclusion
Mat. Mat. LiterQA (L1) 76.47 +0.20 70.85 -12.54 66.90 +4.60 65.31 +1.95 54.92 +16.04 31.50 +28.89
Mat. Comp Extr (L2) 56.16 +28.69 62.07 +24.26 0.00 +79.31 39.53 +24.88 0.12 +73.65 2.54 +49.31
Mat. Data Extr (L2) 88.82 +0.59 88.24 +1.76 90.59 -5.30 82.35 +0.00 61.76 +22.36 70.59 -1.18
Mat. DU (L2) 90.25 +0.25 89.25 -0.25 88.75 -0.25 85.25 +0.75 83.75 +4.75 67.75 +12.50
Mat. Text Sum (L2) 4.82 -0.04 4.60 +0.14 1.18 +3.24 4.55 +0.07 3.30 -1.44 2.66 +0.68
Mat. HV (L2) 90.00 -5.67 68.67 +4.00 65.00 -1.33 64.67 -5.67 66.33 -25.33 19.33 +43.34
Mat. RI (L2) 98.33 -0.28 95.54 +1.40 98.33 -1.12 96.94 +0.00 90.53 +6.96 69.64 +15.32
Val Elec Diff Calc (L3) 46.58 +8.21 46.58 +2.05 28.77 +10.96 19.86 +10.28 23.29 +14.38 23.29 +6.85
Mat. Calc (L3) 37.93 +0.86 35.06 -6.90 27.59 +1.43 19.54 +7.18 20.69 +6.03 38.51 -11.50
Latt Vol Calc (L3) 53.12 -4.37 39.37 -8.12 50.00 +0.62 41.25 +3.75 5.63 +49.37 5.63 +29.37
Perov. Stab Pred (L3) 61.88 -0.63 35.63 +0.20 35.42 +8.12 34.17 +7.29 25.62 +17.71 23.54 +3.33
Diff Rate Analys (L3) 69.13 +26.84 69.80 -1.34 38.93 +15.43 24.83 -0.67 16.11 +28.86 21.48 +4.69
Mat. SafetyQA (L4) 86.21 +1.07 83.35 -12.96 80.02 +2.26 78.72 +2.61 71.70 +6.90 33.41 +37.34
Mat. Tox Pred (L4) 67.15 -1.95 66.50 -3.09 63.90 +3.90 50.08 +7.48 30.41 +37.23 20.65 +37.56
Prop Usage Analysis (L5) 2.68 -0.53 2.75 -0.49 1.08 +0.31 1.15 +0.66 1.97 -0.53 1.46 -0.36
Cry Struct Comp Analys (L5) 1.29 +0.05 1.13 +0.13 1.12 +0.11 1.03 -0.01 1.02 +0.19 1.01 +0.10
Spec Band Gap Gen (L5) 2.51 +0.32 2.70 +0.14 1.00 +1.04 1.28 -0.03 1.08 +0.29 1.01 +0.61
Phys. Phys. LiterQA (L1) 87.56 -0.71 78.91 -8.39 77.61 +2.47 72.96 +1.72 63.58 +17.25 35.97 +31.28
Fund Phys. Exam (L1) 96.59 +0.13 90.48 -6.69 89.26 +3.46 86.44 +1.31 81.01 +10.15 29.43 +56.30
Phys. DU (L2) 99.50 +0.25 99.25 -1.00 99.25 +0.00 98.25 -0.75 97.00 +2.00 80.00 +15.75
Phys. Text Sum (L2) 4.88 -0.08 4.79 -0.01 1.02 +3.51 4.69 +0.00 2.99 -0.80 2.66 +0.85
Phys. HV (L2) 98.75 +0.50 96.00 +1.25 95.25 +0.75 91.75 +0.75 95.00 -2.75 16.00 +70.50
Phys. RI (L2) 99.75 +0.25 98.75 -0.75 99.50 +0.00 99.00 -0.75 97.00 +2.75 90.25 +7.25
HS Phys. Calc (L3) 57.45 -3.44 37.11 -1.58 30.09 +1.86 34.96 +8.02 33.52 +1.87 12.32 +19.06
Gen Phys. Calc (L3) 39.62 -1.00 33.37 -3.87 34.75 +5.50 23.87 +7.26 23.25 +11.75 25.87 -3.24
Phys. Formula Deriv (L3) 4.86 -0.35 4.57 -0.03 1.22 +2.18 1.00 +2.20 1.00 +2.22 1.38 +0.68
Phys. SafetyQA (L4) 88.01 -1.17 81.29 -11.41 81.29 0.87 76.32 3.50 69.59 +10.53 31.58 +38.89
Phys. Lab Safety Test (L4) 78.71 +2.45 75.41 -7.15 70.30 +0.11 67.49 +4.08 62.71 +8.03 20.96 +38.21
Phys. Prob Solving (L5) 3.86 -0.16 2.64 +0.23 1.01 +0.54 1.02 +0.36 1.08 +0.27 1.05 +0.05
A1.3 Detailed Performance of LLMs on Each Task
@@
Table A3: Zero-shot performance of LLMs on each task in the biology domain.
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 M21 M22 M23 M24 M25 M26 Bio LiterQA (L1) 0.8415 0.8371 0.8006 0.8160 0.8050 0.7938 0.7644 0.7667 0.8151 0.8047 0.7716 0.7466 0.7482 0.7206 0.4386 0.7136 0.6400 0.4985 0.7187 0.6746 0.7282 0.7294 0.6165 0.6096 0.6058 0.3980 Prot Prop Iden. (L1) 0.3700 0.3420 0.3173 0.2740 0.4093 0.3013 0.2793 0.4013 0.4587 0.2500 0.2827 0.3407 0.2560 0.2640 0.2520 0.2627 0.2627 0.1687 0.3693 0.2540 0.2053 0.2480 0.2220 0.2093 0.2053 0.1753 Protein Cap. (L1) 0.1298 0.1219 0.1103 0.0477 0.0745 0.1080 0.1180 0.0693 0.1103 0.1140 0.1002 0.1081 0.1120 0.1004 0.0967 0.0156 0.1096 0.1160 0.0906 0.0737 0.0091 0.0171 0.0652 0.0567 0.0333 0.0616 Drug-Drug RE (L2) 0.1307 0.1741 0.0973 0.1687 0.1227 0.1704 0.1602 0.1536 0.1555 0.1720 0.1455 0.1256 0.1675 0.1693 0.0110 0.1024 0.1304 0.0923 0.1488 0.1606 0.1486 0.1104 0.0647 0.1141 0.1017 0.1318 Bio JI (L2) 0.9852 0.9546 0.9134 0.9483 0.9324 0.8712 0.9620 0.9187 0.9388 0.9176 0.9155 0.8701 0.8912 0.8733 0.9060 0.0116 0.7054 0.8110 0.8448 0.9641 0.9271 0.8817 0.2608 0.9261 0.8501 0.5871 C-D RE (L2) 0.1691 0.3770 0.2955 0.3117 0.1372 0.3130 0.3585 0.2370 0.3477 0.3669 0.2888 0.2161 0.2890 0.2127 0.0452 0.2613 0.1906 0.2662 0.2293 0.2292 0.5026 0.2107 0.1334 0.1080 0.2416 0.0902 G-D RE (L2) 0.3929 0.3620 0.2359 0.2382 0.0230 0.4000 0.2918 0.2757 0.3197 0.4859 0.0779 0.0535 0.3916 0.0874 0.0078 0.2514 0.1355 0.3685 0.1234 0.2740 0.1038 0.3066 0.1265 0.1151 0.0743 0.1255 Bio DU (L2) 0.9952 0.9940 0.9952 0.9928 0.9915 0.9915 0.9867 0.9771 0.9940 0.9928 0.9831 0.9879 0.9879 0.9758 0.1775 0.9710 0.9396 0.9312 0.9831 0.9879 0.9650 0.9408 0.9601 0.9287 0.8237 0.7186 Bio Text Summ. (L2) 4.8104 4.7701 4.8104 4.3197 4.8118 4.7709 4.7670 4.7291 4.7709 4.7941 4.8599 4.5967 4.5944 4.6200 4.0495 3.6176 4.2353 4.3963 4.3676 4.5875 2.9574 1.9133 2.8947 3.6246 1.2765 2.5834 Bio HV (L2) 0.9547 0.9482 0.9676 0.9450 0.9385 0.9450 0.9498 0.8900 0.9385 0.9417 0.9094 0.8932 0.9045 0.8560 0.8689 0.1424 0.8074 0.8722 0.8835 0.9061 0.8948 0.6084 0.6197 0.8204 0.5777 0.4223 Bio RI (L2) 0.9815 0.9769 0.9784 0.9753 0.9815 0.9769 0.9660 0.9552 0.9799 0.9784 0.9630 0.9645 0.9676 0.9522 0.3858 0.9599 0.8966 0.9336 0.9460 0.9552 0.9321 0.9090 0.9167 0.9090 0.7855 0.7978 Solu. Pred (L3) 0.4686 0.4831 0.5121 0.5266 0.5024 0.4976 0.4686 0.4686 0.5459 0.5121 0.5169 0.4879 0.4928 0.4928 0.5604 0.5459 0.4734 0.4734 0.4686 0.5217 0.5362 0.5700 0.4686 0.4879 0.4831 0.1836 $\beta$-LA Pred (L3) 0.5369 0.5025 0.5123 0.5123 0.5025 0.4926 0.4384 0.4778 0.4975 0.5025 0.4877 0.4926 0.5025 0.2759 0.5025 0.4483 0.5025 0.0000 0.5025 0.5025 0.4778 0.4975 0.6158 0.4975 0.4975 0.0197 Fluo. Pred (L3) 0.4975 0.5172 0.4384 0.5074 0.5025 0.4631 0.5172 0.5025 0.4975 0.5025 0.5025 0.5222 0.4975 0.4631 0.4975 0.1872 0.5123 0.4286 0.4975 0.4975 0.4483 0.5025 0.2069 0.5025 0.4828 0.0049 GB1 Pred (L3) 0.2823 0.3254 0.2919 0.2919 0.1675 0.3014 0.2440 0.2057 0.2440 0.2440 0.1962 0.3445 0.1914 0.1340 0.2249 0.1435 0.2440 0.0909 0.3062 0.2679 0.2823 0.2488 0.1627 0.2201 0.2201 0.1914 Stab. Pred (L3) 0.3088 0.2402 0.1471 0.2059 0.2892 0.2500 0.3039 0.2647 0.1520 0.2745 0.2696 0.2598 0.2451 0.2255 0.2206 0.1176 0.2304 0.2500 0.2500 0.2549 0.2892 0.2549 0.2647 0.2353 0.2745 0.3725 Prot-Prot Inter. (L3) 0.3140 0.3527 0.3333 0.2609 0.3043 0.2174 0.2754 0.2560 0.2415 0.1981 0.1498 0.1498 0.2126 0.2222 0.1546 0.1159 0.2222 0.1884 0.2222 0.2367 0.2560 0.2609 0.2077 0.2560 0.2415 0.2415 Bio Cal. (L3) 0.5833 0.5833 0.5500 0.5667 0.4167 0.3833 0.3833 0.3167 0.5000 0.5000 0.4167 0.3333 0.3333 0.3667 0.3167 0.3500 0.3167 0.2167 0.2667 0.2500 0.4167 0.2667 0.2167 0.2833 0.2333 0.3167 Bio HarmfulQA (L4) 0.9933 0.5556 0.8721 1.0000 0.9091 0.9192 1.0000 0.9764 0.9327 0.9596 0.5488 0.4916 0.9832 0.5488 0.8687 0.3266 0.7542 0.9865 0.7475 0.0337 0.1515 0.0000 0.4343 0.0135 0.0034 0.0000 Proteotox. Pred (L4) 0.8235 0.8588 0.7941 0.7765 0.6314 0.7902 0.4088 0.4706 0.7529 0.7686 0.5569 0.3971 0.3824 0.3971 0.0618 0.3735 0.3147 0.3735 0.3794 0.3353 0.3706 0.2588 0.2206 0.3088 0.2382 0.0853 Bio Safe Test (L4) 0.8454 0.8667 0.7667 0.7444 0.8299 0.8351 0.7111 0.7111 0.8557 0.7423 0.8402 0.6667 0.6444 0.6778 0.4889 0.5889 0.6667 0.4667 0.6444 0.5667 0.6556 0.5778 0.6111 0.5222 0.3556 0.2889 Bio Proc. Gen (L5) 3.0711 3.1334 2.6638 2.7851 2.6944 3.0087 2.8614 2.1698 2.7782 2.5078 2.2010 2.3102 2.1906 2.2340 2.2582 1.0289 1.5303 1.8839 1.8856 2.0312 1.1317 1.0260 1.0855 1.1478 1.0017 1.1361 Bio Reag. Gen (L5) 2.5556 2.5333 2.3418 2.3929 2.2222 2.3863 2.3707 2.1060 2.2427 2.4000 1.9573 2.2171 2.2393 2.0884 1.9609 1.0359 1.4813 1.9026 1.6564 1.6003 1.0272 1.1077 1.0940 1.1721 1.0120 1.0000 Protein Des. (L5) 0.0670 0.0486 0.0454 0.0316 0.0374 0.0448 0.0437 0.0374 0.0379 0.0325 0.0237 0.0223 0.0263 0.0151 0.0278 0.0525 0.1684 0.0180 0.0143 0.0199 0.0394 0.0162 0.0139 0.0137 0.0110 0.0215 Cell Ana. (L5) 0.0197 0.0165 0.0156 0.0038 0.0007 0.0168 0.0183 0.0054 0.0074 0.0138 0.0014 0.0073 0.0006 0.0033 0.0000 0.0015 0.0038 0.0043 0.0032 0.0021 0.0014 0.0018 0.0008 0.0000 0.0004 0.0007